OpenStudy (anonymous):

Find the slope and the y-intercept of the equation y − 3(x − 1) = 0.

OpenStudy (anonymous):

for y intercept put x=0 in the equation

OpenStudy (shiraz14):

y − 3(x − 1) = 0 y = 3(x-1) = 3x - 3 Gradient of slope = 3; y-intercept = -3